#13
13) A random sample of 50 voters found that 46% were going to vote for a certain 13 candidate. Find the 95% limit for the population proportion of voters who will vote for that candidate.
Answer #1

Sample proportion = 0.46
Sample standard error = SQRT(0.46 x (1 - 0.46)/50) = 0.07
Z-value for 95% CI = 1.96
So, margin of error = 1.96 x 0.07 = 0.137
So, the 95% limit for the population proportion of voters who will vote for the candidate: (0.46 - 0.137 , 0.46 + 0.137) = (0.323 , 0.597)